OpenStudy (anonymous):

(5+5-2x2)/6 Please help I can't figure this out

OpenStudy (anonymous):

5+5=10 2X2=4 10-4=6 (5+5-2x2)/6 (10-4)/6 6/6 =1

OpenStudy (jenniferkeryakous):

I think you just do 5+5 then do 2x2 then minus them and divide by 6 if thats what you mean by /6.

OpenStudy (anonymous):

yes.